Five tables and eight chairs cost $115; three tables and five chairs cost $70. Determine the cost of each table and each chair.: Five tables and eight chairs cost $115; three tables and five chairs cost $70. Determine the cost of each table and each chair.
Answer by oscargut(667)

Let:
x= cost of a chair
y= cost of a table
Five tables and eight chairs cost $115; then 5y+8x=115
three tables and five chairs cost $70. then 3y+5x=70
then:
8x+5y=115
5x+3y=70
1st eq by 5
2nd eq by -8
40x+25y=575
-40x-24y=-560
adding
y=15
5x+3(15)=70
then
5x+45=70
5x=25
x=5
Answer: x=5 and y=15